Operations Report 8-Sept-2021

MDRS Complete Operations Report 08-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 62.9

Minimum 29

Maximum 100

Note on solar: Not sure why this happened. 29 is very low. The only new power use was plugging in the rovers and that happened late in the day. I will keep an eye on it today.

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %

Propane Reading, director tank— 75%

Propane Reading, intern tank— 72%

Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons (and it looks like we are missing a couple of gas cans)

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 350 gallons

Static to Loft Pump used – yes

Water in GreenHab – 140 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 200 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes

Hours: 217.8

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: yes

Hours: 147.2

Beginning charge: 100

Ending charge: 87

Currently charging: no

Spirit rover used: yes

Hours: 135.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: yes

Hours: 72.3

Beginning charge: 100

Ending charge: 95

Currently charging: no

Curiosity rover used: yes

Hours: 141.3

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: The remaining four rovers, with their new batteries, were delivered to the entrance to Cow Dung Road, at 1 pm. We returned them to the station. Opportunity overheated (a past issue) and so was finally retrieved after dark. Spirit’s ignition needs adjusting. Otherwise all are in fine condition, except that they need a conditioner on their exterior surfaces and Percy needs tags.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): All three 350’s

Reason for use: geology and transport to end of road to wait for rovers

Oil Added? No

ATV Fuel Used: 5 Gals

# Hours the ATVs were Used today: 3

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, hauling water

CrewCar used and why, where? Yes, but the spare tire has come loose and we need to fix it before we use the car again

Luna used and why, where? Yes, to town and to retrieve rovers

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Radios and earpieces have been tested

Campus wide inspection, if action taken, what and why? Work party projects in this area are completed

Summary of general operations: It was a slower day. We have been working so hard and were all exhausted.

Summary of Hab operations: Lift boom and power washer were delivered for the paint work. We plan to focus on this large job for the remainder of the work party.

Summary of Outpost operations: General organization and cleanup of area was done. The door to the generator shack was repaired and rehung! Work party projects in this area are completed

Summary of GreenHab operations: General organization and cleanup of GreenHab was done. Work party projects in this area are completed.

Summary of ScienceDome operations: Lift boom and power washer were delivered for the paint work. We plan to focus on this large job for the remainder of the work party.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Still too hot.

Questions, concerns, supplies needed and requests: Can someone investigate and find out what kind of conditioner we should use on the rover’s exterior plastic parts? They are fading and I would like them not to flake. Thank you.